### v2

@@ 3.3 Possible treatment failure: low viral diffusion
In the previous situation, viral particles can reach the outer region of the tumour without any obstacle and the constraints of cell movement do not stop the propagation of the infection. We now analyse the behaviours that arise for small $D_{v}$ and focus on the influence of cell displacement to the spread of the infection.
Figure 4: Comparison in one spatial dimension between the various models with a low value of viral diffusion $D_{v}=10^{-5}\;$mm2/h at time t=1500t=1500\;h. Panels (a-b) show numerical simulations of the discrete model with undirected cell movement (solid lines), the numerical solution of Eq. (U2) (dashed lines) and the numerical solution of Eq. (U3) (dotted black lines) with the parameters given in Table 1, low viral diffusion and both combinations of $\alpha$ and $q_{v}$. Panel (c) show numerical simulations of the discrete model with pressure-driven cell movement (solid lines), the numerical solution of Eq. (P2) (dashed lines) and the numerical solution of Eq. (P3) (dotted black lines) with the parameters given in Table 1, low viral diffusion, $\alpha=3500\;$viruses/cells and $q_{v}=1\;$h-1; we also consider $R_{v}=R_{u}$ in the initial conditions to model a wide initial infection. All the graphical elements have the same meaning as in Fig. 2.
@@
Lower viral diffusion clearly implies a slower invasion speed of infection, as also shown in Fig. 1. The movement of cells is still sufficient for spreading the infection all over the tumour, as shown in Fig. 4a-b. It is also interesting to observe that lower values of $\alpha$ and $q_{v}$ are now associated with a slower speed, differently from the case of faster viral diffusion: indeed, viral particles with a shorter lifespan cannot effectively propagate for long distances. Fig. 4a also shows that some stochastic events slow down the infection to the left of the domain, even though the results are averaged over five simulations; it appears therefore that stochasticity becomes more relevant for this less effective infection, although the overall behaviour is not affected significantly. Fig. 4b shows that for higher values of $\alpha$ and $q_{v}$ the behaviour is almost indistinguishable to the one of Eq. (U2): this is indeed the parameter range in which the quasi-steady assumption of the virus is most accurate. In both cases, the theoretical predictions of the peak of uninfected cells is very precise for the PDE and still constitutes a good approximation for the agent-based model.
@@
As the spatial spread of viral particles becomes less relevant, we expect to recover the main features of Eq. (P2). First of all, a central infection remains mostly confined at the centre of the tumour for both parameter sets that we consider (only shown in the electronic supplementary material S4): the central infection causes the total cell density to drop, while external uninfected cells proliferate; since cells cannot move towards an area with higher cell density and viral diffusion is negligible, the outer cells are never going to be infected and the tumour keeps expanding in the same way it would do in the absence of treatment, almost identically to the situation described in Ref. [morselli23].
@@
We therefore focus on initial infections spread in the whole tumour, namely $R_{v}=R_{u}$, which gives rise to the formation of travelling waves for Eq. (P2) [morselli23]. For the sake of brevity, we restrict our attention to higher values of $\alpha$ and $q_{v}$, which facilitate the spread of the infection for the continuous model; we remark that both cases are shown in electronic supplementary material S5. Fig. 4c shows that the same kind of travelling waves is also obtained as a solution of Eq. (P3) and the two solutions superimpose almost exactly; the peak of uninfected cells is higher than the one of the solutions of Eqs. (U2) and (U3), in line with the less effective spread of the infection. Note that the agent-based model again shows that the infection is confined to the centre of the tumour, in line with the case that neglects explicit viral dynamics. It is important to stress that, for the current model, demographic stochasticity plays a much more important role than in previous models: any growth above the average of uninfected cells stops the movement of infected cells and hence cannot be compensated at later times by other processes.
@@
Taking the average over a higher number of realisations does not significantly improve the agreement, since the individual stochasticity of each simulation is not affected (not shown). According to the formal derivation of the PDEs from the agent-based model, an increase of cell number, ensuring that the evolution of the stochastic system diverges less significantly from the average behaviour described by the continuum model, and a decrease of the temporal and spatial discretisation all togheter improve the quality of the continuum approximation. In this specific case, the PDE predicts the presence of a very small infected cell density up to the uninfected invasion front; a good agreement would therefore require a number of infected cells high enough to neglect stochasticity. However, such an agreement can only be obtained at scales that are much larger than any meaningful biological process.
@@
Figure 5: Numerical simulation of the discrete model with pressure-driven movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, with the exception of the diffusion coefficient of viral particles $D_{v}$ (which is set to $10^{-5}\;$mm2/h), the death rate of infected cells $q$ (which is set to $8.33\times 10^{-3}\;$h-1, i.e. one-fifth of the reference values) and the initial radius of viral infection $R_{v}$ (which is set equal to $R_{u}$); furthermore, we use the values $\alpha=3500\;$viruses/cells and $q_{v}=1\;$h-1. The dashed cyan circles in the top panels represent the expected positions of the uninfected invasion fronts in the absence of treatment, travelling at speed $\sqrt{D_{u}p/2}$. The dashed red circles in the bottom panels represent the front of the infected cells given by the numerical solution of Eq. (P3).
@@
Reasonable increases of the effective infection rate $\tilde{\beta}$ do not lead to a better infiltration of the infection, as this causes a decrease in central cell density and creates the need for infected cells to move against a pressure gradient. Also, as we will show shortly, oscillations emerge in some parameter ranges associated with an effective infection, but are suppressed at low values of viral diffusion. On the other hand, building upon the results of Ref. [morselli23], we expect that decreasing the death rate of infected cells $q$ leads to a better outcome.
@@ 3.4 Significant treatment success: emergence of oscillations
As we anticipated, it is well known that solutions of Eq. (U3) may exhibit persistent oscillations for some parameter values; this is in line with the fact that the non-spatial model of Eq. (2.6) presents a Hopf bifurcation. Fig. 6a shows a codimension one bifurcation plot in $\alpha$, whereas Fig. 6b-c shows the time series for three different values of $\alpha$. We remark that a similar behaviour is also observed, for example, as $\beta$} increases or as $q_{v}$ decreases. The one-dimensional setting is not ideal for studying oscillations, as local tumour extinctions due to stochasticity are very common: indeed, the typical behaviour observed in such parameter ranges is analogous to the one described below in relation to Fig. 10. We therefore concentrate our attention to two spatial dimensions: the higher number of cells involved allow us to describe oscillations, as well as stochastic-driven partial and complete eradication.
@@
Figure 6: (a) One parameter bifurcation plot in $\alpha$ of Eq. (2.2), with other parameters as in Table 1 and $q_{v}=0.167\,$h-1. The green dots show the maximum and minimum values of $u$ for the stable limit cycle oscillations. The solid lines show the equilibrium value of $u$; the line is red if the equilibrium is stable and black if it is unstable. Hopf bifurcations are denoted by HB. Observe that for low values of $\alpha$, the infection-free equilibrium close to carrying capacity is stable. (b-c) Numerical simulation of Eq. (2.5) with the parameters as in Table 1 and different values of the viral burst size $\alpha$. The solid blue lines represent uninfected cells, the solid red lines infected cells and the dotted purple lines represent viral density multiplied by $q_{v}/(\alpha q)$; the two latter quantities superimpose almost exactly. The oscillations become wider as $\alpha$ increases, in accordance with the bifurcation diagram of panel (a).
@@
Figure 7: Numerical simulations of the discrete model with undirected movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1. Panel (a) shows uninfected cell density from a single simulation to show the impact of stochasticity; the dashed cyan circle represents the expected positions of the tumour invasion front, travelling at speed $2\sqrt{Dp}$. Panels (b) and (c) represent cell densities on the horizontal section of the domain $L,L]\times\{0\}$, obtained by averaging five simulations: the density of the uninfected tumour cells is the blue solid line and the density of the infected tumour cells is the red solid line (not shown in panel (b) to avoid the superimposition with uninfected cell density); the viral density is not shown to improve readability. The solutions of Eq. (U3) are represented with dotted black lines. The vertical black dashed lines represent the expected positions of the uninfected invasion front, travelling at speed $2\sqrt{Dp}$. The horizontal solid black lines show the equilibrium of the ODE given by Eq. (2.7) and the horizontal dashed yellow line represents the expected uninfected density at the front given by Eq. (3.2) (only relevant in panel (c)). The maximum of the axes and the colorbars were scaled to enhance readability and are much lower than the maximum over time of the quantity plotted.
@@
Fig. 7, along with the video accompanying it (see electronic supplementary material S7), shows an overall excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in two spatial dimensions. After the initial infection reaches the tumour boundary, the number of cancer cells decreases significantly; as a consequence, stochasticity becomes relevant and, in some regions, cells go extinct. Fig. 7a shows the inhomogeneous pattern of uninfected cells caused by this. Similar patterns would also arise in the two-dimensional non-radial continuum model under stochastic perturbations. As time elapses, the spatial inhomogeneities quickly disappear, and both models show similar persistent oscillations in the tumour centre around the equilibrium values (coherently with what the bifurcation diagram shows). A slight delay in the oscillations is the main difference between the two modelling approaches: this is easy to see from the total number of cells, depicted in Fig. 9a. The invasion speed of the tumour in the surrounding area is still $2\sqrt{Dp}$ and the uninfected cell density at the front is approximately given by Eq. (3.2), despite some fluctuations. It is interesting to note that highly dynamic patterns may emerge when viral injections at multiple sites are considered [baabdulla23]: this can be easily checked through VisualPDE [walker23], using the simulation available at the link https://visualpde.com/sim/?mini=raZybpXG.
Fig. 9a shows the comparison with the solution of Eq. (U2) to remark once again that oscillations cannot be obtained in the two-equations model even for high infection rates. Furthermore, Fig. 9b shows that oscillations are almost non-existent for lower viral diffusion, resembling the behaviour of Eq. (U2); this confirms once again the excellent quality of the quasi-steady assumption as spatial viral dynamics become negligible and the reactions involving the virus increase their rate.
Figure 8: Numerical simulations of the discrete model with pressure-driven movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1. All the graphical elements have the same meaning as per Fig. 7, with the vertical black dashed lines representing the expected positions of an untreated tumour, travelling at speed $\sqrt{Dp/2}$ and the vertical black dash-dotted lines representing the expected invasion speed in presence of virotherapy, given by Eq. (A.3). Observe that in this second case we consider a wave starting from an initial radius of 33\;$m$ rather than $R_{u}=2.6\;$$m$, in order to compensate the transient behaviour in which the infection is far from the boundary of the tumour.
@@
Both models have an invasion speed much lower than $\sqrt{Dp/2}$: indeed, the nonlinear diffusion is affected by the scale of the cell density, as explained in Appendix A. In our case, the oscillatory behaviour of the system does not allow to easily identify a unique, invariant speed. Fig. 8 shows that we can obtain a good approximation by setting $\bar{u}=1800\;$cells/mm2, which corresponds to an average of the peak over time. The change in the speed in presence of effective infections is a significant difference from the case of undirected cell movement and, as a result, the overall cell numbers involved are lower in the second case (see Fig. 9a). It is also interesting to remark that inhomogeneous infection patterns may result in an anisotropic spread of the tumour in the surrounding environment: we refer the interested reader to the VisualPDE simulation available at the link https://visualpde.com/sim/?mini=e0LGK0st.
Let us remark once again that oscillations are only possible when the viral diffusion is high enough. If this is not the case, the behaviour of Eq. (P3) is well approximated by Eq. (P2), central infections remain constrained to the centre of the tumour and wide infections lead to a travelling wave without relevant oscillations, and the agent-based model results into a localised infection in all the relevant parameter ranges, as shown in Fig. 9c.
@@
The oscillations analysed so far are able to drive the tumour at low densities, in which stochastic fluctuations may cause the death of all the agents in some regions; nevertheless, oscillations are not wide enough to lead the whole tumour close to extinction. It is important to see if it is possible to increase the effectiveness of the therapy and have the possibility of a full tumour remission. We therefore decrease the value of $q_{v}$ to half of the reference value, so that the oscillations can reach lower cell densities. We also consider an initial wide viral infection (i.e., $R_{v}=R_{u}$) to study the effect of a higher initial viral load. Fig. 10a-b summarise the cell number of one hundred simulations for both cell movement at time t=1500t=1500\;h. In both cases, the wide majority of the simulations shows the extinction of the tumour; the case of undirected cell movement appears more successful, with only eight simulations with a surviving tumour (in contrast to nineteen in the other case) and an overall lower cell density at the final time. It is convenient to track the distribution over time of the extinctions of the agent-based model through the tumour control probability, defined as
@@
Figure 10: (a)-(b) Violin plots of the final cell number at time t=1500t=1500\;h obtained from one hundred simulations of the agent-based models with the parameters given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=8.33\times 10^{-2}\;$h-1; the latter has been set to the half of the lowest reference values to model a more effective viral infection. We also consider $R_{v}=R_{u}$ in the initial conditions to model a wide initial infection. The blue dots show the single results greater than zero. The white dots show the median, which is zero in both cases. (c)–(d) Tumour and infection control probabilities for the two agent-based models from Eqs. (3.3) and Eq. (3.4) (undirected movement in solid blue and pressure-driven movement in solid red) and their corresponding continuous counterparts via the Poissonian assumption from Eqs. (3.5) and Eq. (3.6) (same colours, with dotted lines).
@@
$\text{TCP}(t)\coloneqq\exp\Bigl[-2\pi\int_{0}^{R}u(t,r)\,\mathrm{d}r\Bigr]$ (3.5)
which is based on the assumption that the number of surviving uninfected cells follows a Poisson distribution. In Ref. [gong13], the Poissonian TCP defined as above was compared with more complex expressions and it was found to perform equally well in the case of radiotherapy. Fig. 10c shows an excellent quantitative agreement at early times between the expressions of Eqs. (3.3) and (3.5). The agreement is clearly lost as the continuum model shows the recrudescence of the tumour. It is also interesting to observe that the maximum value of the Poissonian TCP is slightly lower than the asymptotic value of the discrete TCP in both cases. A possible explanation for this discrepancy is that localised cell clusters are more effectively infected than wider, spatially homogeneous populations with a lower density; in the agent-based models, stochastic fluctuations during the infection naturally generate such small clusters, whereas the continuum model represents only averages and therefore cannot reproduce this effect. Nevertheless, the Poissonian TCP still provides a meaningful insight on the underlying stochastic model.
@@
$\text{ICP}(t)\coloneqq\exp\Bigl[-2\pi\int_{0}^{R}i(t,r)\,\mathrm{d}r\Bigr]$ (3.6)
@@
Our results show that, at least in principle, the virotherapy may cause a complete tumour remission if the viral infection is extremely effective; this is interesting, but it remains an open question how to translate this in a clinical settings. We remark that, for such low cell densities, the immune clearance of the surviving cells could facilitate tumour extinction; on the other hand, an early immune response may hinder the infection and lead to an overall decrease of the therapy effectiveness, as discussed in Ref. [morselli24].

### v3

@@ 3.3 Possible treatment failure: low viral diffusion
In the previous situation, viral particles can reach the outer region of the tumour without any obstacles and the constraints of cell movement do not stop the propagation of the infection. We now analyse the behaviours that arise for small $D_{v}$ and focus on the influence of cell displacement to the spread of the infection. We remark that the case of fast viral dynamics and slow diffusion is the one in which we expect the quasi-steady assumption to be extremely accurate; as a consequence, we also expect Eqs. (U2) and (P2) to approximate well Eqs. (U3) and (P3).
Figure 4: Comparison in one spatial dimension between the various models with a low value of viral diffusion $D_{v}=10^{-5}\;$mm2/h at time t=1500t=1500\;h. Panels (a-b) show numerical simulations of the discrete model with undirected cell movement (solid lines), the numerical solution of Eq. (U2) (dashed lines) and the numerical solution of Eq. (U3) (dotted black lines) with the parameters given in Table 1, low viral diffusion and both combinations of $\alpha$ and $q_{v}$. Panel (c) shows numerical simulations of the discrete model with pressure-driven cell movement (solid lines), the numerical solution of Eq. (P2) (dashed lines) and the numerical solution of Eq. (P3) (dotted black lines) with the parameters given in Table 1, low viral diffusion, $\alpha=3500\;$viruses/cells and $q_{v}=1\;$h-1; we also consider $R_{v}=R_{u}$ in the initial conditions to model a wide initial infection. All the graphical elements have the same meaning as in Fig. 2.
@@
Lower viral diffusion clearly implies a slower invasion speed of infection, as also shown in Figs. 1 and 12b. The movement of cells is still sufficient for spreading the infection all over the tumour, as shown in Fig. 4a-b. It is also interesting to observe that lower values of $\alpha$ and $q_{v}$ are now associated with a slower speed, differently from the case of faster viral diffusion: indeed, viral particles with a shorter lifespan cannot effectively propagate for long distances. Fig. 4a also shows that some stochastic events slow down the infection to the left of the domain, even though the results are averaged over five simulations; it appears therefore that stochasticity becomes more relevant for this less effective infection, although the overall behaviour is not affected significantly. Fig. 4b shows that for higher values of $\alpha$ and $q_{v}$ the behaviour is almost indistinguishable to the one of Eq. (U2): this is indeed the parameter range in which the quasi-steady assumption of the virus is most accurate. In both cases, the theoretical predictions of the peak of uninfected cells is very precise for the PDE and still constitutes a good approximation for the agent-based model.
@@
As the spatial spread of viral particles becomes less relevant, we expect to recover the main features of Eq. (P2). First of all, a central infection remains mostly confined at the centre of the tumour for both parameter sets that we consider (only shown in the electronic supplementary material S4): the central infection causes the total cell density to drop, while external uninfected cells proliferate; since cells cannot move towards an area with higher cell density and viral diffusion is negligible, the outer cells are never going to be infected and the tumour keeps expanding in the same way it would do in the absence of treatment, almost identically to the situation described in Ref. [53].
@@
We therefore focus on initial infections spread in the whole tumour, namely $R_{v}=R_{u}$, which gives rise to the formation of travelling waves for Eq. (P2) [53]. For the sake of brevity, we restrict our attention to higher values of $\alpha$ and $q_{v}$, which facilitate the spread of the infection for the continuous model; we remark that both cases are shown in electronic supplementary material S5. Fig. 4c shows that the same kind of travelling waves is also obtained as a solution of Eq. (P3) and the two solutions superimpose almost exactly; the peak of uninfected cells is higher than the one of the solutions of Eqs. (U2) and (U3), in line with the less effective spread of the infection. Note that the agent-based model again shows that the infection is confined to the centre of the tumour, in line with the case that neglects explicit viral dynamics. It is important to stress that, for the current model, demographic stochasticity plays a much more important role than in previous models: any growth above the average of uninfected cells stops the movement of infected cells and hence cannot be compensated at later times by other processes.
@@
Taking the average over a higher number of realisations does not significantly improve the agreement, since the individual stochasticity of each simulation is not affected (not shown). According to the formal derivation of the PDEs from the agent-based model, an increase of cell number, ensuring that the evolution of the stochastic system diverges less significantly from the average behaviour described by the continuum model, and a decrease of the temporal and spatial discretisation all together improve the quality of the continuum approximation. In this specific case, the PDE predicts the presence of a very small infected cell density up to the uninfected invasion front; a good agreement would therefore require a number of infected cells high enough to neglect stochasticity. However, such an agreement can only be obtained at scales that are much larger than any meaningful biological process.
@@
Figure 5: Numerical simulation of the discrete model with pressure-driven movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, with the exception of the diffusion coefficient of viral particles $D_{v}$ (which is set to $10^{-5}\;$mm2/h), the death rate of infected cells $q$ (which is set to $8.33\times 10^{-3}\;$h-1, i.e. one-fifth of the reference values) and the initial radius of viral infection $R_{v}$ (which is set equal to $R_{u}$); furthermore, we use the values $\alpha=3500\;$viruses/cells and $q_{v}=1\;$h-1. The dashed cyan circles in the top panels represent the expected positions of the uninfected invasion fronts in the absence of treatment, travelling at speed $\sqrt{D_{P}p/2}$. The dashed red circles in the bottom panels represent the front of the infected cells given by the numerical solution of Eq. (P3).
@@
Reasonable increases of the effective infection rate $\tilde{\beta}$ do not lead to a better infiltration of the infection, as this causes a decrease in central cell density and creates the need for infected cells to move against a pressure gradient. Also, as we will show shortly, oscillations emerge in some parameter ranges associated with an effective infection, but are suppressed at low values of viral diffusion. On the other hand, building upon the results of Ref. [53], we expect that decreasing the death rate of infected cells $q$ leads to a better outcome.
@@ 3.4 Significant treatment success: emergence of oscillations
As we anticipated, it is well known that solutions of Eq. (U3) may exhibit persistent oscillations for some parameter values; this is in line with the fact that the non-spatial model of Eq. (2.6) presents a Hopf bifurcation, as rigorously analysed in Refs. [4, 6]. Fig. 6a shows a codimension one bifurcation plot in $\alpha$, whereas Fig. 6b-c shows the time series for three different values of $\alpha$. We remark that a similar behaviour is also observed, for example, as $\beta$} increases or as $q_{v}$ decreases. The one-dimensional setting is not ideal for studying oscillations, as local tumour extinctions due to stochasticity are very common: indeed, the typical behaviour observed in such parameter ranges is analogous to the one described below in relation to Fig. 10. We therefore concentrate our attention to two spatial dimensions: the higher number of cells involved allows us to describe oscillations, as well as stochastic-driven partial and complete eradication.
@@
Figure 6: (a) One parameter bifurcation plot in $\alpha$ of Eq. (2.2), with other parameters as in Table 1 and $q_{v}=0.167\,$h-1. The green dots show the maximum and minimum values of $u$ for the stable limit cycle oscillations. The solid lines show the equilibrium value of $u$; the line is red if the equilibrium is stable and black if it is unstable. The Hopf bifurcation is denoted by HB. Observe that for low values of $\alpha$, the infection-free equilibrium close to carrying capacity is stable. (b-c) Numerical simulation of Eq. (2.5) with the parameters as in Table 1 and different values of the viral burst size $\alpha$. The solid blue lines represent uninfected cells, the solid red lines infected cells and the dotted purple lines represent viral density multiplied by $q_{v}/(\alpha q)$; the two latter quantities superimpose almost exactly. The oscillations become wider as $\alpha$ increases, in accordance with the bifurcation diagram of panel (a).
@@
Figure 7: Numerical simulations of the discrete model with undirected movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1. Panel (a) shows uninfected cell density from a single simulation to show the impact of stochasticity; the dashed cyan circle represents the expected positions of the tumour invasion front, travelling at speed $2\sqrt{D_{U}p}$. Panels (b) and (c) represent cell densities on the horizontal section of the domain $L,L]\times\{0\}$, obtained by averaging five simulations: the density of the uninfected tumour cells is the blue solid line and the density of the infected tumour cells is the red solid line (not shown in panel (b) to avoid the superimposition with uninfected cell density); the viral density is not shown to improve readability. The solutions of Eq. (U3) are represented with dotted black lines. The vertical black dashed lines represent the expected positions of the uninfected invasion front, travelling at speed $2\sqrt{D_{U}p}$. The horizontal solid black lines show the equilibrium of the ODE given by Eq. (2.7) and the horizontal dashed yellow line represents the expected uninfected density at the front given by Eq. (3.2) (only relevant in panel (c)). The maximum of the axes and the colorbars were scaled to enhance readability and are much lower than the maximum over time of the quantity plotted.
@@
Fig. 7, along with the video accompanying it (see electronic supplementary material S7), shows an overall excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in two spatial dimensions. After the initial infection reaches the tumour boundary, the number of cancer cells decreases significantly; as a consequence, stochasticity becomes relevant and, in some regions, cells go extinct. Fig. 7a shows the inhomogeneous pattern of uninfected cells caused by this. Similar patterns would also arise in the two-dimensional non-radial continuum model under stochastic perturbations. As time elapses, the spatial inhomogeneities quickly disappear, and both models show similar persistent oscillations in the tumour centre around the equilibrium values (coherently with what the bifurcation diagram shows). A slight delay in the oscillations is the main difference between the two modelling approaches: this is easy to see from the total number of cells, depicted in Fig. 9a. The invasion speed of the tumour in the surrounding area is still $2\sqrt{D_{U}p}$ and the uninfected cell density at the front is approximately given by Eq. (3.2), despite some fluctuations. It is interesting to note that highly dynamic patterns may emerge when viral injections at multiple sites are considered [4]: this can be easily checked through VisualPDE [74], using the simulation available at the link https://visualpde.com/sim/?mini=raZybpXG.
Fig. 9a shows the comparison with the solution of Eq. (U2) to remark once again that oscillations cannot be obtained in the two-equations model even for high infection rates. Furthermore, Fig. 9b illustrates that oscillations are almost non-existent for lower viral diffusion, resembling the behaviour of Eq. (U2); this confirms once again the excellent quality of the quasi-steady assumption as spatial viral dynamics become negligible and the reactions involving the virus increase their rate.
Figure 8: Numerical simulations of the discrete model with pressure-driven movement in two spatial dimensions at three different times. The parameters employed are the ones given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1. All the graphical elements have the same meaning as per Fig. 7, with the vertical black dashed lines representing the expected positions of an untreated tumour, travelling at speed $\sqrt{D_{P}p/2}$ and the vertical black dash-dotted lines representing the expected invasion speed in presence of virotherapy, given by Eq. (A.3). Observe that in this second case we consider a wave starting from an initial radius of 33\;$m$ rather than $R_{u}=2.6\;$$m$, in order to compensate the transient behaviour in which the infection is far from the boundary of the tumour.
@@
Both models have an invasion speed much lower than $\sqrt{D_{P}p/2}$: indeed, the nonlinear diffusion is affected by the scale of the cell density, as explained in Appendix A. In our case, the oscillatory behaviour of the system does not allow to easily identify a unique, invariant speed. Fig. 8 shows that we can obtain a good approximation by setting $\bar{u}=1800\;$cells/mm2, which corresponds to an average of the peak over time. The change in the speed in presence of effective infections is a significant difference from the case of undirected cell movement and, as a result, the overall cell numbers involved are lower in the second case (see Fig. 9a). It is also interesting to remark that inhomogeneous infection patterns may result in an anisotropic spread of the tumour in the surrounding environment: we refer the interested reader to the VisualPDE simulation available at the link https://visualpde.com/sim/?mini=e0LGK0st.
Let us remark once again that oscillations are only possible when the viral diffusion is high enough. If this is not the case, the behaviour of Eq. (P3) is well approximated by Eq. (P2), central infections remain constrained to the centre of the tumour and wide infections lead to a travelling wave without relevant oscillations. The agent-based model results into a localised infection in all the relevant parameter ranges, as shown in Fig. 9c.
@@
The oscillations analysed so far are able to drive the tumour to low densities, where stochastic fluctuations may cause the death of all the agents in some regions; nevertheless, oscillations are not wide enough to lead the whole tumour close to extinction. It is important to see if it is possible to increase the effectiveness of the therapy and have the possibility of a full tumour remission. We therefore decrease the value of $q_{v}$ to half of the reference value, so that the oscillations can reach lower cell densities. We also consider an initial wide viral infection (i.e., $R_{v}=R_{u}$) to study the effect of a higher initial viral load. Fig. 10a-b summarise the cell numbers of one hundred simulations for both cell movement at time t=1500t=1500\;h. In both cases, the wide majority of the simulations shows the extinction of the tumour; the case of undirected cell movement appears more successful, with only eight simulations with a surviving tumour (in contrast to nineteen in the other case) and an overall lower cell density at the final time. It is convenient to track the distribution over time of the extinctions of the agent-based model through the tumour control probability, defined as
@@
Figure 10: (a)-(b) Violin plots of the final cell number at time t=1500t=1500\;h obtained from one hundred simulations of the agent-based models with the parameters given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=8.33\times 10^{-2}\;$h-1; the latter has been set to half of the lowest reference value to model a more effective viral infection. We also consider $R_{v}=R_{u}$ in the initial conditions to model a wide initial infection. The blue dots show the single results greater than zero. The white dots show the median, which is zero in both cases. (c)–(d) Tumour and infection control probabilities for the two agent-based models from Eqs. (3.3) and Eq. (3.4) (undirected movement in solid blue and pressure-driven movement in solid red) and their corresponding continuous counterparts via the Poissonian assumption from Eqs. (3.5) and Eq. (3.6) (same colours, with dotted lines).
@@
$\text{TCP}(t)\coloneqq\exp\Bigl[-2\pi\int_{0}^{R}u(t,r)r\,\mathrm{d}r\Bigr]$ (3.5)
which is based on the assumption that the number of surviving uninfected cells follows a Poisson distribution. In Ref. [29], the Poissonian TCP defined as above was compared with more complex expressions and it was found to perform equally well in the case of radiotherapy. Fig. 10c shows an excellent quantitative agreement at early times between the expressions of Eqs. (3.3) and (3.5). The agreement is clearly lost as the continuum model shows the recrudescence of the tumour. It is also interesting to observe that the maximum value of the Poissonian TCP is slightly lower than the asymptotic value of the discrete TCP in both cases. A possible explanation for this discrepancy is that localised cell clusters are more effectively infected than wider, spatially homogeneous populations with a lower density; in the agent-based models, stochastic fluctuations during the infection naturally generate such small clusters, whereas the continuum model represents only averages and therefore cannot reproduce this effect. Nevertheless, the Poissonian TCP still provides a meaningful insight on the underlying stochastic model.
@@
$\text{ICP}(t)\coloneqq\exp\Bigl[-2\pi\int_{0}^{R}i(t,r)r\,\mathrm{d}r\Bigr]$ (3.6)
@@
Building upon the strong quantitative agreement observed so far, we now focus on the continuum model and examine how parameter variations influence the extinction probability. We adopt the same parameter settings as in the simulations shown in Fig. 10 and investigate how the viral burst size $\alpha$ affects the maximal value attained for positive times by $\text{TCP}(t)$ and $\text{ICP}(t)$ in the numerical solutions of Eqs. (U3) and (P3). Note that the maximal value of $\text{TCP}(t)$ corresponds to the probability of tumour extinction within the time window considered, since in the agent-based model no new cancer cells can arise after extinction has occurred. The situation is different in the case of $\text{ICP}(t)$, as an infection can emerge even in absence of infected cells. In particular, the initial conditions in Eq. (3.1) yield $\text{ICP}(0)=1$, which does not reflect the subsequent dynamics: the initial viral load inevitably triggers infection (see also Fig. 10d). By contrast, extinction of infected cells at later times generally indicates complete eradication of the infection, due to the rapid decay of viral particles. For this reason, we restrict our analysis to $t>0$ when computing the maximal values of both probabilities over time.
Fig. 11 shows that the tumour extinction has a higher probability as $\alpha$ increases, confirming once again that a higher viral burst is associated to a more effective treatment. For $\alpha$ high enough the probability is very close to one, meaning that in practice we can always expect tumour extinction; such values however are probably beyond the biologically meaningful range. It is also interesting to observe that the eradication of the infection is much more common and could be associated either to the eradication of the full tumour (for high values of $\alpha$) or to the total failure of the treatment (for low values of $\alpha$).
Our results show that, at least in principle, virotherapy may cause a complete tumour remission if the viral infection is extremely effective; this is interesting, but it remains an open question how to translate it in a clinical settings. We remark that, for such low cell densities, the immune clearance of the surviving cells could facilitate tumour extinction; on the other hand, an early immune response may hinder the infection and lead to an overall decrease of the therapy effectiveness, as discussed in Ref. [54].
Figure 11: Maximal values attained at positive time by $\text{TCP}(t)$ and $\text{ICP}(t)$ for the numerical solutions of Eqs. (U3) and (P3) for different values of $\alpha$. The other parameters are the ones listed in Table 1, with the exception of $q_{v}$ that is set to $8.33\times 10^{-2}\;$h-1 (i.e., half of the lowest reference value) to model a more effective viral infection. We also consider $R_{v}=R_{u}$ in the initial conditions to model a wide initial infection.